How many moles of NH3 are produced when 2.1 mol of H2 reacts?
1
Expert's answer
2020-06-19T06:59:58-0400

Solution.

2,1 x

3H2 + N2 = 2NH3

3 2

x mol NH3; 2,1 mol H2.

x = 2,1*2/3 = 1,4 mol.

x = n(NH3) = 1,4 mol.

Answer: 1,4 mol NH3